OpenStudy (anonymous):

multiply (x^7+x^6)(x^7+x^6-x^5)

OpenStudy (anonymous):

(x^7+x^6)(x^7+x^6-x^5) = x^7(x^7+x^6-x^5) + x^6 (x^7+x^6-x^5) = x^(7+7) +x^(7+6) - x^(7+5) +x^(6+7) + x^(6+6) -x^(5+6) =x^14 +x^13 -x^12 +x^13 + x^12 - x^11 =x^14 + 2x^13 -x^11